OpenStudy (breanna.rae):

is anyone one in Connections academy?? Algerbra 1B Lesson 8 Unit 1

OpenStudy (tennis5518):

yes

OpenStudy (tennis5518):

im in flvs

OpenStudy (breanna.rae):

Im doing a Work Sample and i need help

OpenStudy (breanna.rae):

Solve the system by substitution. 2x+2y=38 y=x+3

OpenStudy (tennis5518):

what do you think?

OpenStudy (breanna.rae):

i honestly have no idea

OpenStudy (tennis5518):

ok wait

OpenStudy (tennis5518):

what are the step to solve do you know

OpenStudy (breanna.rae):

i dont :(

OpenStudy (tennis5518):

ok so first you need to Divide thru 1st by 2 x + y = 19

OpenStudy (tennis5518):

Next you need to Substitute for "y" and solve for "x": x + x+3 = 19 2x = 16

OpenStudy (tennis5518):

After x = 8 -------- solve for "y": x + y = 19 8 + y = 19 y = 11

OpenStudy (tennis5518):

so what the answer i told you every thing!

OpenStudy (breanna.rae):

Thank you so much!

OpenStudy (tennis5518):

what the answer can you explain it now?

OpenStudy (breanna.rae):

the answer is 11, right?

OpenStudy (tennis5518):

yes can you say it all again

OpenStudy (tennis5518):

say it all again with out copy and paste

OpenStudy (breanna.rae):

what do you mean, say it all again

OpenStudy (tennis5518):

what did i tell you can you explain

OpenStudy (tennis5518):

also medal and fan also testimonal plz

OpenStudy (breanna.rae):

divide the first equation 2x+2y=38 38/2=19 x+y=19 and im lost at what you did after that

OpenStudy (breanna.rae):

@Tennis5518

OpenStudy (tennis5518):

niceeee good job! now close this question!

OpenStudy (breanna.rae):

okay thank you again

OpenStudy (anonymous):

anyone do the unit 2 lesson 9 algebra b unit test?